OpenStudy (anonymous):

What is 3 to the power of 2 over 3 equal to? cube root of 9 square root of 9 cube root of 27 square root of 27

OpenStudy (cj49):

square root of 9

OpenStudy (anonymous):

oh you ment square root of 9?

OpenStudy (jdoe0001):

\(\Large \bf { \\ \quad \\ a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}} \\ \quad \\ \quad \\ 3^{\frac{{\color{blue}{ 2}}}{{\color{red}{ 3}}}}\implies ? }\)

OpenStudy (cj49):

\[3^{2}\div3\]

OpenStudy (anonymous):

that would be 9 /3 which is 2

OpenStudy (cj49):

its 3x3=9

OpenStudy (cj49):

3x3/3=3

OpenStudy (solomonzelman):

yeah... it is not so exactly written. It can be as jdoe001 said 2\(\scriptsize\color{slate}{^{\LARGE ⅔} }\), or it can be 2²/3.

OpenStudy (jdoe0001):

*cough* it helps to use parentheses to group the proper ones *cough*

OpenStudy (solomonzelman):

yeah the options sound more like jdoe

OpenStudy (solomonzelman):

yeah parenthesis, or a "\(\scriptsize\color{slate}{^{\LARGE ⅔} }\)" symbol

OpenStudy (jdoe0001):

\(\Large \bf { a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}} \\ \quad \\ \quad \\ 3^{\frac{{\color{blue}{ 2}}}{{\color{red}{ 3}}}}\implies \sqrt[\square ?]{\square ?} }\)

OpenStudy (anonymous):

@jdoe0001sooo is it \[\sqrt[3]{3^2}\]

OpenStudy (anonymous):

@jdoe0001 or wait no it would be \[\sqrt[2]{3^3}\]

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\large \sqrt[2]{3^3}\implies \sqrt{27}\)

OpenStudy (anonymous):

so now i simplify and its the last one right

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks a bunch ;)

OpenStudy (solomonzelman):

No, this is not right. The original question is \(\normalsize\color{black}{ 3^{\bf\LARGE⅔}}\). \(\normalsize\color{black}{ 3^{\bf\LARGE⅔}=\sqrt[3]{3^2}=\sqrt[3]{9}}\)

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

ohhh ahemm yea.... right... hold the mayo and onions =)

OpenStudy (jdoe0001):

the 1st one you had was correct, myu bad =) so

OpenStudy (jdoe0001):

\(\Large \bf { a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}} \\ \quad \\ \quad \\ 3^{\frac{{\color{blue}{ 2}}}{{\color{red}{ 3}}}}\implies \sqrt[{\color{red}{ 3}}]{3^{\color{blue}{ 2}}}\implies \sqrt[3]{9} }\)

OpenStudy (solomonzelman):

yeah... nothing wrong calculations wise, just confused what the problem was. That is not bad at all.

OpenStudy (anonymous):

oh okayas right the first time with the first problem i did before i corrected it lol so the first one?